OpenStudy (anonymous):

6 + 3x - x3 intervals of increase. So far I got the critical numbers to be 1 and -1. I think this increases from -infin to -1. and decreases from -1,1. then increases again from 1 to infinity. ' Is this right?

OpenStudy (anonymous):

I think you have to differentiate and equate it to zero or something!

jhonyy9 (jhonyy9):

- so you need to check it for : - x=-1 --- 6-3(-1)-1*3=6+3-1=8 x=-2 --- 6-6-8=-8 x=0 --- 6+0-0 =6 x=1 --- 6+3-1=8 - so you can see that is true exactly how you have wrote it.